# Supervised Learning-enhanced Multi-Group Actor Critic for Live-stream Recommendation

### v1

@@ 6. Conclusion
In this work, we propose a novel Supervised Learning-enhanced Multi-Group Actor-Critic algorithm (SL-MGAC) to optimize request-level live-stream injection policies in the challenging context of a live-stream & short video mixed recommendation system. Unlike existing RL methods for recommendation systems, our approach focuses on enhancing the stability and robustness of the RL model. Specifically, we introduce a multi-group state decomposition module to reduce prediction variance and seamlessly integrate multi-task reward learning with traditional critic learning to constrain Q-value estimation. In practice, we aim to minimize the risk of policy deterioration or model collapse, thereby enabling the proposed SL-MGAC method to be successfully deployed in large-scale industrial recommendation systems.
7. Acknowledgment
We acknowledge Bai Wei and Chen Xiaoshuang for proposing detailed modification advice to help us improve the quality of this manuscript.
Appendix A Proof of Theorem 3.1
Proof.
Let state space $\mathcal{S}=\bigcup_{i=1}^{K}\mathcal{S}_{i},\forall i\neq j\;\mathcal{S}_{i}\cap\mathcal{S}_{j}=\varnothing$, state $s\in\mathcal{S}$. Let $r$ be the real reward of state $s$ and action $a$, $\hat{r}$ be the predicted value of $r$ over $\mathcal{S}$, $\hat{r}_{i}$ be the predicted reward of $r$ over $\mathcal{S}_{i}$.
Since a user has at most one request at timestamp $t$, a user will have at most one state $s$. We assume the state count $N$ of $\mathcal{S}$ is finite, the state count of $\mathcal{S}_{i}$ is $N_{i}$. Let $\mu$ be the mean value of $\hat{r}$ over $\mathcal{S}$, $\mu_{i}$ be the mean value of $\hat{r}_{i}$ over $\mathcal{S}_{i}$, then we have:
(16) $Var[\hat{r}]$ $=\frac{1}{N}\sum_{j=1}^{N}\left(\hat{r}_{j}-\mu\right)^{2}$
$=\frac{1}{N}\sum_{i=1}^{K}\sum_{j=1}^{N_{i}}\left(\hat{r}_{ij}-\mu\right)^{2}$
$=\frac{1}{N}\sum_{i=1}^{K}\sum_{j=1}^{N_{i}}\left(\hat{r}_{ij}-\mu_{i}+\mu_{i}-\mu\right)^{2}$
$=\frac{1}{N}\sum_{i=1}^{K}\sum_{j=1}^{N_{i}}\left(\hat{r}_{ij}-\mu_{i}\right)^{2}+\frac{1}{N}\sum_{i=1}^{K}N_{i}\left(\mu_{i}-\mu\right)^{2}$
where the first term is within-class variance $Var_{a}[\hat{r}]$ and the second term is between-class variance $Var_{b}[\hat{r}]$. As $Var_{b}[\hat{r}]>0$, we have
(17) $Var[\hat{r}]>\frac{1}{N}\sum_{i=1}^{K}\sum_{j=1}^{N_{i}}\left(\hat{r}_{ij}-\mu_{i}\right)^{2}=Var_{a}[\hat{r}]\$
As we partition the state space $\mathcal{S}$ into $K$ disjoint sub-spaces $\mathcal{S}_{i}$, then the variance of a neural network $f$ with multi-group state decomposition(MG SD) module is $Var_{a}[\hat{r}]$ while the variance of a neural network $f^{\prime}$ without MG-SD module is $Var[\hat{r}]$. ∎
Appendix B Tables
@@
We denote MGN as the multi-group network, RPN as the Reward Prediction Network, QRN as Q Residual Networks, DDB as Distribution Discretization Bins. Both live and video watch time distribution have 7 time bins. We leave a separate time bin for the $a=1$ case and hence the output layer dimension of RPN is 8.
Appendix C Figures
Figure 8. Training Curves of SL-MGAC and SL-MGAC-sep.
Figure 9. Adaptive adjustments of live-stream exposure during peak hours and low-traffic periods.
Figure 10. The amplitude distribution of live-stream injection ratio.

### v2

@@ 7. Conclusion
In this work, we propose a novel Supervised Learning-enhanced Multi-Group Actor-Critic algorithm (SL-MGAC) to optimize request-level live stream injection policies in the challenging context of a live stream & short video mixed recommendation system. Unlike existing RL methods for recommendation systems, our approach focuses on enhancing the stability and robustness of the RL model. Specifically, we introduce a multi-group state decomposition module to reduce prediction variance and seamlessly integrate multi-task reward learning with traditional critic learning to constrain Q-value estimation. In practice, we aim to minimize the risk of policy deterioration or model collapse, thereby enabling the proposed SL-MGAC method to be successfully deployed in large-scale industrial recommendation systems.
Appendix A Tables
@@
We denote MGN as the multi-group network, RPN as the Reward Prediction Network, QRN as Q Residual Networks, and DDB as Distribution Discretization Bins. Both live and video watch time distribution have 7 time bins. We leave a separate time bin for the $a=1$ case, and hence the output layer dimension of RPN is 8.
Appendix B Figures
Figure 7. Data distributions of online live stream room (or viewer) count, live stream watch time and short video watch time.
Figure 8. The amplitude distribution of live stream injection ratio.